You want 350 liters of a solution with 12% acid. You have two solutions, one is 16% and the other is 9%. How many liters of the 16% solution will be needed?
Found 2 solutions by Alan3354, ankor@dixie-net.com:
Answer by Alan3354(69443)

Do it like this one. Only the numbers are different.
-----------
Amanda wants to make 8 gal. of a 20% saline solution by mixing together a 56% saline solution and a 8% saline solution. How much of each solution must she use?
================
e = amount of 8%
f = amount of 56%
---
e+ f = 8 (total solution)
8e + 56f = 20*8 (total saline)
---
e+ f = 8
e + 7f = 20
------------ Subtract
-6f = -12
f = 2
e = 6

Answer by ankor@dixie-net.com(22740)

You want 350 liters of a solution with 12% acid. You have two solutions, one is 16% and the other is 9%.
How many liters of the 16% solution will be needed?
:
Let x = amt of 16% solution
total is to be 350 liters, therefore
(350-x) = amt of 9% solution
:
.16x + .09(350-x) = .12(350)
.16x + 31.5 - .09x = 42
.16x - .09x = 42 - 31.5
.07x = 10.5
x = 10.5/.07
x = 150 liters of the 16% solution